OpenStudy (anonymous):

Please Helppp!!! Want a medal If g(x) = 2(x - 4), find the value of x if g(x) = 20. 32 12 14 10

jigglypuff314 (jigglypuff314):

solve for x for 20 = 2(x - 4)

OpenStudy (sheraz12345):

14

OpenStudy (anonymous):

why is it 14 ? can you explain

OpenStudy (anonymous):

oh did not see your post jiggly puff

OpenStudy (sheraz12345):

its damn easy.. Put 20=2(X-4) MULTIPLY 2 WITH X-4 AND SIMPLIFY

jigglypuff314 (jigglypuff314):

solve for x for 20 = 2(x-4) so first distribute you would get 20 = 2x - 8 then add 8 to both sides +8 +8 28 = 2x then divide both sides by 2 /2 /2

OpenStudy (anonymous):

thanks jiggly puff

OpenStudy (anonymous):

ignore my pic please !!! LOL

jigglypuff314 (jigglypuff314):

lol glad I could help :)